OpenStudy (anonymous):

Express entire radical as mixed radical: √180

jimthompson5910 (jim_thompson5910):

Similar Problem: \[\Large \sqrt{18} = \sqrt{9*2}\] \[\Large \sqrt{18} = \sqrt{9}*\sqrt{2}\] \[\Large \sqrt{18} = \sqrt{3^{2}}*\sqrt{2}\] \[\Large \sqrt{18} = 3\sqrt{2}\]